LawnTech sells its sonic gopher repeller at a price of $380, and estimates price elasticity of demand at -2.6 at this price. At that price, LawnTech sells 35,000 units per year. LawnTech is thinking about cutting its price by -3.5%. How many MORE units can LawnTech expect to sell at the new price? Rounding: tenth of a unit.
Answer #1

E = (% change in quantity demanded/% change in price)

-2.6 = % change in quantity demanded/(-3.5%)

% change in quantity demanded = -2.6*(-3.5%) = 9.1%

So,

More units to be sold = 35000*9.1%

More units to be sold = 3185 units